OpenStudy (anonymous):

Is there a number that is exactly 4 more than its cube?

OpenStudy (saifoo.khan):

x - x^3 = 4

OpenStudy (anonymous):

Well, lets call such a number k. That would imply that \[k^3 + 4 = k\]\[\implies k^3 - k + 4 = 0\] So now we just need to find the roots of this polynomial.

OpenStudy (anonymous):

It doesn't look like it'll be a terribly pretty number, but since it's an odd degree polynomial it must have at least one root, so the number does exist.

OpenStudy (anonymous):

It won't. All three roots are irrational.

OpenStudy (anonymous):

We can find the exact solutions for this equations. They involve cube roots of quadratic surds.